OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation 5(2x + 3) = 85 is given below. Choose the step that is missing from the solution. 5(2x + 3) = 85 10x = 70 x = 7

OpenStudy (anonymous):

Are 10x = 70 and x = 7 the answer options?

OpenStudy (anonymous):

10x + 3 = 85 2x + 15 = 85 10x + 15 = 17 10x + 15 = 85

OpenStudy (anonymous):

those are the options

OpenStudy (anonymous):

You are using the associative property as the next mathematical procedure/step. 5(2x+3) = 10x + 15 which should be on the left hand side of the equation that should go 2nd.

OpenStudy (anonymous):

so C?

OpenStudy (anonymous):

D